# An Information Theoretic Treatment of Animal Movement ${}^{\dagger}$

### v1

@@ 2 Segmentation Hierarchy
$*{}^{*}$ $(\sigma_{i})$ $i=1,\cdots,n$ $\mu f$ secs $n$ clusters of segments and $\mu$ steps
@@
${}^{{\dagger}}$ $(\kappa_{c})$ $c=1,\cdots,k$ $m\mu f$ secs $k$ clusters of words and $m$ StaMEs
@@
${}^{{\ddagger}}$ ${}^{**}$ sub-diel homogeneous: string of same CAMs
@@
[…]
age & ${}^{{\ddagger}{\ddagger}}$ seasonal internal and external factors
${}^{\parallel}$ syndromic individual’s LiMP specific DARs condition on
[…]
$*{}^{*}$tatistical movement elements, formerly called metaFuMEs
${}^{{\dagger}}$anonical activity mode segments, formerly called “short duration” or homogeneous CAMs
${}^{{\ddagger}}$ehavioral activity mode segments, formerly called “long duration” or heterogeneous CAMs
[…]
${}^{\parallel}$ife-time track, including dispersal, migration, and nurturing young
${}^{**}$o be determined using biological change point analysis [6, 7] and hidden Markov methods [12, 14, 15]
${}^{{\dagger}{\dagger}}$ay be extracted using appropriate clustering methods [24]
${}^{{\ddagger}{\ddagger}}$o be determined using appropriate statistical methods [26, 27]
[…]
@@
$\mathcal{T}^{\sigma}$ (StaME track) $(i;\sigma_{i})$ $\lfloor\frac{T}{\mu}\rfloor H(\mathbf{p}^{\sigma})$ $*{}^{*}$ 1st level centroids
@@
[…]
${}^{+}$he difference between raw ($\kappa_{c}$) and rectified ($\kappa_{c}^{\star}$) CAMs based on Eq 3.14 (see Fig 1)
$*{}^{*}$nformation encoded into $\mathcal{T}^{\rm loc}$ if symbol sequences are uncorrelated
${}^{{\dagger}}$ivergence across the ensemble distributions $\mathbf{p}^{\mathcal{W}_{c}}$, $c=1,\cdots,k$ (Eq 5)
${}^{\ddagger}$vg. divergence of $\omega_{\ell}$ distributions in $\mathcal{W}_{c}$ from their distribution in $\mathcal{W}$ itself
[…]
The current concept of a CAM also needs to be treated more rigorously than before if it is to be integrated into an information theoretic framework. Specifically, in the formulation below, $m$ $\mu$-length segments are string together into $m\mu$-step words (Task 3; Fig 1, Tables 1 and 2) that then constitute a word set $\mathcal{W}$. The words in $\mathcal{W}$ can then be used to generate a word time series $\mathcal{T}^{\rm wd}$ of $\lfloor T/(m\mu)\rfloor$ elements (Table 2). A cluster analysis of the words in $\mathcal{W}$ is then used to identify $k$ word clusters denoted by $\mathcal{W}_{c}$, $c=1,\ldots,k.$ The centroids of these $k$ clusters can be used to generate the $k$-element raw CAM set $\mathcal{K}$ (Task 4, Fig 1), with elements $\kappa_{c}$ used to code track $\mathcal{T}^{\rm wd}$ into a raw CAM track $\mathcal{T}^{\kappa}$ (Table 2).
Additionally, one may also translate each word into its representative $m$-symbol string and then identify which cluster most words that have the same $m$-symbol code belong to. If this is the word set $\mathcal{W}_{c}$, then all words with this particular $m$-symbol coding sequence will be identified with a rectified CAM type $c$, as represented by its archetype/ideal $\kappa^{\star}_{c}$ (Task 5, Fig 1). The corresponding rectified CAM coding track is denoted by $\mathcal{T}^{\rm CAM}$. The proportion $E^{\mathcal{\kappa}}$ of these particular $m$-symbol words that had to be reassigned, when added for all $c=1\cdots,k$, provides a measure of the reliability of the CAM coding process. In a nutshell, the difference between the $\mathcal{T}^{\kappa}$ and $\mathcal{T}^{\rm CAM}$ time series is that the former arises from a cluster analysis of segments containing $m\mu$ steps and the latter assigns a CAM type $\kappa^{\star}_{c}$ to all words that are coded by the same $m$-symbol sequence of StaMEs $\sigma_{i}\in\mathcal{S}$. This distinction will be made more rigorous in the next section where we formalize the set of Tasks 1-6 (Fig 1) that define the functor $\mathcal{M}:\mathcal{T}^{\rm loc}\mapsto\left\{\mathcal{T}^{\sigma},\mathcal{T}^{\rm CAM},E^{\mathcal{\kappa}}\right\}$ used to derive the StaME and rectified CAM sets and track codings, and the assignment error associated with the process.
@@ 3 Two Level Track Segmentation
[…]
@@ 5 Information Content and Method Comparisons
[…]
$\mathcal{M}_{a}$ ::: ${\mathcal{T}}_{a}\mapsto\{\mathcal{S}_{a},\mathcal{K}^{\star}_{a},E^{\mathcal{\kappa}_{a}}\},\quad a=1,2,\cdots,$
@@
Beyond using the entropy measures $H^{\mathcal{K}^{\star}_{a}}$ (Eq 4.4) and misassignment error rates $E^{\mathcal{\kappa}_{a}}$ (Eq 3.16) of different functors $a$ to compare various approach and identify those with high coding and low error rates, we can compare the CAM distributions ${\bf p}^{\kappa^{\star}}$ in sets $\mathcal{K}^{\star}_{a}$ that have the same number $k$ of CAMs using the information theoretic Kullback-Leibler and Jensen-Shannon divergence measures [37]. The Kullback-Leibler divergence measure allows us to compute the “distances” (divergences) between two distributions. Specifically, if $\mathbf{p}^{1}=\left(p^{1}_{1},\cdots,p^{1}_{k}\right)$ and $\mathbf{p}^{2}=\left(p^{2}_{1},\cdots,p^{2}_{k}\right)$ are two different distributions then the Kullback-Leibler measure of the divergence of distribution $\mathbf{p}^{2}$ from distribution $\mathbf{p}^{1}$ is
@@
$=$ $H\big{(}{\mathbf{p}}^{\rm mix}\big{)}-\sum_{c=1}^{k}p^{\kappa}_{c}H\big{(}{\mathbf{p}}^{\mathcal{W}_{c}}\big{)}$

### v2

# An Information Theory Treatment of Animal Movement Tracks

@@ 2 Segmentation Hierarchy
StaME $(\sigma_{i})$ $i=1,\cdots,n$ $\mu f$ secs $n$ clusters of segments and $\mu$ steps
@@
CAM $(\kappa_{c})$ $c=1,\cdots,k$ $m\mu f$ secs $k$ clusters of words and $m$ StaMEs
@@
BAM TBD∗∗ sub-diel homogeneous: string of same CAMs
@@
LiMP depends on multi-day to LiMP specific DARs condition on
age & env.‡‡ seasonal internal and external factors
LiT syndromic individual’s LiMP specific DARs condition on
move. types§§ life span internal and external factors
Statistical movement elements, formerly called metaFuMEs
Canonical activity mode segments, formerly called “short duration” or homogeneous CAMs
Behavioral activity mode segments, formerly called “long duration” or heterogeneous CAMs
§Diel activity routine segments
Life-time movement phase segments, including dispersal, migration, and nurturing young
Life-time track, including dispersal, migration, and nurturing young
∗∗To be determined using biological change point analysis [6, 7] and hidden Markov methods [12, 14, 15]
††May be extracted using appropriate clustering methods [24]
‡‡To be determined using appropriate statistical methods [26, 27]
§§For discussions on movement syndrome types see [28, 29]
@@
$\mathcal{T}^{\sigma}$ (StaME track) $(i;\sigma_{i})$ $\lfloor\frac{T}{\mu}\rfloor H(\mathbf{p}^{\sigma})$ bits 1st level centroids
@@
$\mathcal{T}^{\rm CAM}$ (rectified CAM track) $(j;\kappa_{c}^{\star})^{+}$ $\lfloor\frac{T}{m\mu}\rfloor H(\mathbf{p}^{\rm CAM})$ bits§ argmax $c$ reassignment
+The difference between raw ($\kappa_{c}$) and rectified ($\kappa_{c}^{\star}$) CAMs based on Eq 3.14 (see Fig 1)
Information encoded into $\mathcal{T}^{\rm loc}$ if symbol sequences are uncorrelated
Divergence across the ensemble distributions $\mathbf{p}^{\mathcal{W}_{c}}$, $c=1,\cdots,k$ (Eq 5)
Avg. divergence of $\omega_{\ell}$ distributions in $\mathcal{W}_{c}$ from their distribution in $\mathcal{W}$ itself
§Information that can be coded into $\mathcal{T}^{\rm loc}$ if CAM sequences are uncorrelated
The current concept of a CAM also needs to be treated more rigorously than before if it is to be integrated into an information theoretic framework. Specifically, in the formulation below, $m$ $\mu$-length segments are strung together into $m\mu$-step words (Task 3; Fig 1, Tables 1 and 2) that then constitute a word set $\mathcal{W}$. The words in $\mathcal{W}$ can then be used to generate a word time series $\mathcal{T}^{\rm wd}$ of $\lfloor T/(m\mu)\rfloor$ elements (Table 2). A cluster analysis of the words in $\mathcal{W}$ is then used to identify $k$ word clusters denoted by $\mathcal{W}_{c}$, $c=1,\ldots,k.$ The centroids of these $k$ clusters can be used to generate the $k$-element raw CAM set $\mathcal{K}$ (Task 4, Fig 1), with elements $\kappa_{c}$ used to code track $\mathcal{T}^{\rm wd}$ into a raw CAM track $\mathcal{T}^{\kappa}$ (Table 2).
Additionally, one may also translate each word into its representative $m$-symbol string and then identify which cluster most words that have the same $m$-symbol code belong to. If this is the word set $\mathcal{W}_{c}$, then all words with this particular $m$-symbol coding sequence will be identified with a rectified CAM type $c$, as represented by its archetype/ideal $\kappa^{\star}_{c}$ (Task 5, Fig 1). The corresponding rectified CAM coding track is denoted by $\mathcal{T}^{\rm CAM}$. The proportion $E^{\mathcal{\kappa}}$ of these particular $m$-symbol words that had to be reassigned, when added for all $c=1\cdots,k$, provides a measure of the reliability of the CAM coding process. In a nutshell, the difference between the $\mathcal{T}^{\kappa}$ and $\mathcal{T}^{\rm CAM}$ time series is that the former arises from a cluster analysis of segments containing $m\mu$ steps and the latter assigns a CAM type $\kappa^{\star}_{c}$ to all words that are coded by the same $m$-symbol sequence of StaMEs $\sigma_{i}\in\mathcal{S}$. This distinction will be made more rigorous in the next section where we formalize the set of Tasks 1-6 (Fig 1) that defines the functor $\mathcal{M}:\mathcal{T}^{\rm loc}\mapsto\left\{\mathcal{T}^{\sigma},\mathcal{T}^{\rm CAM},E^{\mathcal{\kappa}}\right\}$ used to derive the StaME and rectified CAM sets and track codings, and the assignment error associated with the process.
@@ 3 Two Level Track Segmentation
    In general, we cannot expect $\mathcal{T}^{\kappa}$ and $\mathcal{T}^{\rm CAM}$ to represent the same sequence of CAMs. Errors will occur and the proportion of errors that occur can be determined by computing the proportion of locations $j$ where the CAM designations in $\mathcal{T}^{\kappa}$ and $\mathcal{T}^{\rm CAM}$ do not match up. This error computation can be either be computed by directly comparing the CAM sequences in $\mathcal{T}^{\kappa}$ and $\mathcal{T}^{\rm CAM}$ or using the proportion of misclassifications that arise in computing the CAM assignment using rule Eq 3.14. In the latter case it is clear that the proportion of times that wordj, when coded as word type $\omega_{\ell}$, is correctly assigned to CAM $\kappa_{c}$ is $\left(\max\limits_{c}\{p^{{\mathcal{W}}_{c}}_{j}\}/\sum_{c=1}^{k}p^{{\mathcal{W}}_{c}}_{j}\right)$: proportional misassigned to another CAM is 1 minus this quantity. Thus, summing over all misassignments of $\omega_{\ell}$, taking into account the frequency $N^{\omega_{\ell}}/N^{\rm wd}$ of these words in $\mathcal{W}$ we obtain the percentage error rate
@@ 5 Information Content and Method Comparisons
Essentially the task of producing the best (among a comparative group of parameter and clustering method approaches) rectified CAMs from relocation data can be expressed in terms of a family of functors $\mathcal{M}_{a}$, $a=0,1,2,\cdots,$. The arguments of these functors are the parameters $\{\mu,n,m,k\}$ and the methods selected to cluster the segments into StaMEs (denoted CM($n$,approachσ)) and the words into CAMs (denoted CM($k$,approachκ)). Following our earlier notation, our family of functors are mappings of a track $\mathcal{T}^{\rm loc}$ onto a set of StaMEs $\mathcal{S}$, rectified CAMs $\mathcal{K}^{\star}$, and error measure $E^{\mathcal{\kappa}}$ all indexed by an approach designator $a$:
$\mathcal{M}_{a}$ ::: $\mathcal{T}^{\rm loc}_{a}\mapsto\{\mathcal{T}^{\sigma}_{a},\mathcal{T}^{\rm CAM}_{a},E^{\mathcal{\kappa}_{a}}\},\quad a=1,2,\cdots,$
@@
Beyond using the entropy measures $H^{\kappa^{\star}_{a}}$ (Eq 4.4) and misassignment error rates $E^{\mathcal{\kappa}_{a}}$ (Eq 3.16) of different functors $a$ to compare various approach and identify those with high coding and low error rates, we can compare the CAM distributions ${\bf p}^{\kappa^{\star}}$ in sets $\mathcal{K}^{\star}_{a}$ that have the same number $k$ of CAMs using the information theoretic Kullback-Leibler and Jensen-Shannon divergence measures [37]. The Kullback-Leibler divergence measure allows us to compute the “distances” (divergences) between two distributions. Specifically, if $\mathbf{p}^{1}=\left(p^{1}_{1},\cdots,p^{1}_{k}\right)$ and $\mathbf{p}^{2}=\left(p^{2}_{1},\cdots,p^{2}_{k}\right)$ are any two discrete distributions with $k$ bins then the Kullback-Leibler measure of the divergence of distribution $\mathbf{p}^{2}$ from distribution $\mathbf{p}^{1}$ is
@@
$=$ $H^{\omega}\big{(}{\mathbf{p}}^{\rm mix}\big{)}-\sum_{c=1}^{k}p^{\kappa}_{c}H^{\omega}\big{(}{\mathbf{p}}^{\mathcal{W}_{c}}\big{)}$